# Vision Language Model-based Testing of Industrial Autonomous Mobile Robots

### v1

@@ Abstract
Autonomous Mobile Robots (AMRs) are deployed in diverse environments (e.g., warehouses, retail spaces, and offices), where they work alongside humans. Given that human behavior can be unpredictable and that AMRs may not have been trained to handle all possible unknown and uncertain behaviors, it is important to test AMRs under a wide range of human interactions to ensure their safe behavior. Moreover, testing in real environments with actual AMRs and humans is often costly, impractical, and potentially hazardous (e.g., it could result in human injury). To this end, we propose a Vision Language Model (VLM)-based testing approach (RVSG) for industrial AMRs developed by PAL Robotics in Spain. Based on the functional and safety requirements, RVSG uses the VLM to generate diverse human behaviors that violate these requirements. We evaluated RVSG with several requirements and navigation routes in a simulator using the latest AMR from PAL Robotics. Our results show that, compared with the baseline, RVSG can effectively generate requirement-violating scenarios. Moreover, RVSG-generated scenarios increase variability in robot behavior, thereby helping reveal their uncertain behaviors.
@@ 2 Industrial Context
Within the RoboSAPIENS, self-adaptive behaviors are implemented using the MAPLE-K loop [8], an extension of the well-established MAPE-K loop for self-adaptation in software systems. As presented in Fig. 1, in the MAPLE-K loop, an AMR continuously monitors (Monitor) its environment using dedicated sensors, analyzes the data (Analyze), and, upon detecting a situation that requires adaptation, plans (Plan), validates (Legitimate), and executes (Execute) the selected adaptation. Throughout this process, all information becomes part of the robot’s knowledge base (Knowledge), which can be accessed at any stage of the loop. The Legitimate phase is newly introduced in MAPLE-K.
@@ 3 Method
3.1 RVSG Overview
Fig. 2 presents an overview of RVSG for generating scenarios during the iterative AMR testing. As shown in Fig. 2, RVSG has two main stages: (i) environment preprocessing and (ii) test scenario generation. In the first stage, RVSG achieves the Map Image Preprocessing and Environment Description Generation. At Map Image Preprocessing step ( 1 in Fig. 2), RVSG first obtains a top-down view of the simulation world as a map image. This map can then be labeled to highlight the spatial structure, facilitating subsequent layout analysis and task planning. Based on the labeled map, RVSG constructs a grid map with precise waypoint mapping to refine the layout and support robot navigation description and human configuration generation. Both processed images are provided to the VLM for Environment Description Generation at 2. Within 2, the VLM utilizes multi-turn conversations to extract key layout and navigation information from the images, and generates a natural language description of the environment, which assists in subsequent test scenario generation.
@@ 3.1 RVSG Overview
Figure 2: Overview of RVSG during the iterative testing of autonomous mobile robots.
For test scenario generation, at 3, RVSG integrates the specific F&S requirement, robot navigation route, and the preprocessed map images and environment description. These inputs are structured into carefully designed prompts (detailed in Fig. 3) for the VLM, guiding its automated understanding and analysis to generate human configurations. At 4, within the Human Configuration Generation by Multi-turn Conversation, Prompt Generator automatically selects appropriate initial multi-turn prompt templates based on the requirement and navigation route, filling them into predefined positions. Similarly, map images and the environment description are incorporated to further enrich the prompt content, providing strong support for VLM reasoning. These prompts are sequentially input into the VLM, enabling step-by-step reasoning through multi-turn conversations to produce intermediate outputs (e.g., scenario description) and final outputs (e.g., JSON format human behavior configuration). Each turn of conversation builds on previous context, ensuring the VLM continuously aligns with the test objective, understands the requirement, and generates the required output for each step.
After completing all conversation turns, RVSG delivers the final formatted output to the Configuration Decoder, which interprets the human configuration generated by the VLM based on the specific requirement and robot navigation route. The dialogue history is stored in the Memory module via Dialogue History Storage for future optimization. Moreover, the Scenario & Human Task Extractor retrieves the scenario description and human tasks (e.g., workers assigned to packing and delivery in a warehouse setting) from the dialogue, and summarizes this information to support experience accumulation for RVSG scenario generation.
The decoded human configuration is used to generate executable human agents in the simulator. To ensure realistic human behavior, RVSG employs HuNavSim [9], an agent controller based on behavior trees and the social force model, which closely simulates real-world human actions. The AMR under test is loaded with its designated navigation route and uses its navigation algorithm for environment perception and navigation decisions. Both human agents and the AMR perform their tasks in the simulator at 5, enabling dynamic performance evaluation of the AMR along the navigation route, with time series data collected throughout the process. The simulation environment is consistent with that used in preprocessing, and real-time communication is maintained via Robot Operating System 2 (ROS2) [10].
The time series data collected during simulation is used by the Objective Function Calculator to compute an objective function value, quantitatively reflecting the evaluation results based on the initial requirement. This result, together with the time series data, is submitted to the Feedback Integrator to generate feedback data. The feedback data can be used to optimize the original prompts in the Human Configuration Generation module for more effective scenario configuration generation. In addition, along with scenario and human task information, it is integrated into the Memory module through the Memory Integrator, which accumulates experience for VLM scenario generation, supporting prompt refinement and increased scenario diversity in future multi-turn conversations.
As shown at 6, with simulation feedback from VLM-generated test scenarios and scenario generation experience stored in the Memory module, the Prompt Generator in Human Configuration Generation can continually refine prompts. This enhances human configuration effectiveness for similar scenarios and increases the diversity of the generated test scenarios, enabling iterative testing. Ultimately, RVSG flexibly produces a more effective and diverse set of test scenarios based on simulation feedback, scenario diversity, and predefined stopping criteria (e.g., generating 10 scenarios), thereby improving the comprehensiveness and effectiveness of AMR testing.
@@ 3.2 Environment Preprocessing
Leveraging the VLM’s strong visual reasoning, RVSG conducts multi-turn conversations to analyze step-by-step the labeled map and grid images, extracting layout and navigation information. As shown in the Environment Description Generation module of Fig. 2, RVSG first provides the labeled map image and layout analysis instructions to the VLM to obtain a detailed analysis of environmental layout. Next, using the preserved dialogue history, RVSG sends the grid image and corresponding grid analysis instructions to the VLM, along with the map image that emphasizes the basic environmental information, and generates descriptions of the grid and navigation layouts. Finally, RVSG provides a summarization prompt to the VLM asking to summarize the key description for both map and grid images in JSON format. This output is then processed by the Environment Description Extractor, which extracts and pairs the corresponding descriptions for each image, storing them together as the current environment description for subsequent testing and analysis. Given the potential for hallucinations or inaccuracies in the VLM’s output, the final environment description is manually reviewed to ensure correctness.
@@ 3.3 Test Scenario Generation
As shown at 3 and 4 in Fig. 2, RVSG utilizes multi-turn conversations to generate scenarios and human configurations based on system requirements, robot navigation routes, and environment descriptions. By incorporating feedback from scenario simulations and memory of previous scenarios (at 6), RVSG further optimizes configuration generation and enhances scenario diversity. Specifically, RVSG provides multiple prompt templates for five key reasoning steps: (1) Test Scenario Description Generation, (2) High-level Human Task Generation, (3) Low-level Human Waypoint Path Generation, (4) Human Configuration Validity Check and Revision, and (5) JSON Format Human Configuration Generation.
Test Scenario Description Generation analyzes the current environment and produces realistic textual descriptions of test scenarios involving human activities. These descriptions serve as the basis for subsequent human configuration. High-level Human Task Generation allocates specific human tasks according to the generated scenario description, detailing concrete workflows and outputting the names of key labeled areas involved in these workflows. Low-level Human Waypoint Path Generation builds on the high-level tasks, sequentially selecting valid waypoints from the preprocessed grid map based on the workflow and key areas to construct human waypoint paths. The high-level human tasks and low-level waypoint paths together form the human configuration, which is then checked for validity by the Human Configuration Validity Check and Revision. Any problematic configurations are automatically detected and revised by the VLM at this stage. For example, the module verifies whether the workflow aligns with the scenario description and whether the waypoint paths are feasible. If inconsistencies are found, the workflow is modified and the waypoint paths are updated accordingly. Finally, JSON Format Human Configuration Generation provides the final human configuration in JSON format.
Fig. 3 presents all seven prompt templates provided by RVSG. Each template focuses on one or two specific key reasoning step(s). Notably, for Test Scenario Description Generation, the module has introduced an additional prompt version (Prompt2 in Fig. 3) beyond the initial prompt (Prompt1), influenced by previous scenario memory. Similarly, to regenerate the human configuration based on simulation feedback without modifying the underlying test scenario, RVSG integrates the prompts for High-level Human Task Generation and Low-level Human Waypoint Path Generation (Prompt3 and Prompt4), and provides a new Human Configuration Regeneration prompt (Prompt5). Based on these templates, RVSG automatically fills in the provided information, such as requirements, into the corresponding fields to generate prompts for VLM reasoning and analysis. For example, Prompt1 in Fig. 3 contains three fields: Environment Description, Robot Navigation Route, and Requirement. RVSG automatically populates each field with relevant content and outputs the final prompt. Note that the content for fields, such as Requirement and Robot Navigation Route, does not adhere to a strict format and can be flexibly adjusted based on experience.
@@
Figure 3: Prompt templates for multi-turn conversations and scenario generation workflow with feedback and memory in RVSG.
@@
RVSG combines different prompts to support iterative testing across various stages. As shown in Fig. 3, during the initial stage, RVSG selects the initial Test Scenario Description Generation prompt (Prompt1), High-level Human Task Generation prompt (Prompt3), Low-level Human Waypoint Path Generation prompt (Prompt4), Human Configuration Validity Check and Revision prompt (Prompt6), and JSON Format Human Configuration Generation prompt (Prompt7). These are combined to form the initial multi-turn conversation prompts that guide the VLM in generating the initial human configuration, which is then executed in the simulator to obtain simulation results. Subsequently, the feedback trigger determines whether to initiate feedback-based optimization. If optimization continues, RVSG combines the dialogue history with Human Configuration Regeneration (Prompt5), Human Configuration Validity Check and Revision (Prompt6), and JSON Format Human Configuration Generation (Prompt7) to guide the VLM in refining the human configuration based on simulation feedback, without altering the initial scenario description. The optimized configuration is then executed in the simulator, after which new feedback is obtained. This feedback includes the objective function value corresponding to the testing requirement, which is used to evaluate the effectiveness of the optimized configuration and determine the degree of optimization. If further optimization is needed, the generated feedback can be used iteratively in subsequent optimization steps. If the feedback trigger indicates stopping feedback optimization (e.g., reaching a certain number of optimizations), RVSG then relies on the memory trigger to decide whether to use memory for generating more diverse scenarios. If memory-based optimization proceeds, RVSG replaces the initial Test Scenario Description Generation prompt (Prompt1) with the Diversity Test Scenario Description Generation prompt (Prompt2), keeping the other prompts unchanged to guide the VLM in creating the new test scenario. Moreover, in Prompt2, the Historical Scenario Memory field provides the descriptions of all previously generated test scenarios, along with their corresponding optimal configurations and objective function values, serving as references for generating more diverse scenarios. Otherwise, the process terminates, and all generated scenarios are output. All multi-turn conversation histories, simulation results, and the corresponding scenario descriptions and human tasks are stored in the Memory module, supporting VLM-based scenario generation at different stages. Note that, throughout all multi-turn conversations, RVSG consistently provides labeled map and grid images either directly or indirectly. For simplicity, these details are omitted in the workflow shown in Fig. 3.
@@ 4.1 Subject System and Simulator
We test the latest logistics AMR robot, TIAGo OMNI Base from PAL Robotics [11]. It has four mecanum wheels for omnidirectional movement and two LiDAR scanners that provide a 360° unobstructed field of view, enabling it to detect and autonomously avoid obstacles to ensure safe navigation. To support simulation and testing, PAL Robotics offers an open-source TIAGO OMNI Base ROS 2 Simulation [12]. The simulator is built on Gazebo classic 11 [13] and provides a high-fidelity model of the TIAGo OMNI Base. The TIAGo OMNI Base adopts Nav2 [14] for both global navigation and local obstacle avoidance, while real-time communication is achieved through ROS 2 Humble [10]. Considering the simulation platform and the ROS 2 version, we selected the compatible HuNavSim [9] as the default agent for human behavior control. At the group level, HuNavSim uses the social force model [15] to simulate general crowd flow and avoidance behaviors in the presence of obstacles or other humans. At the individual level, it employs behavior trees to define and control more complex, personalized human actions. As a result, HuNavSim can provide realistic human behavior modeling, which benefits AMR navigation task testing and supports the development of navigation systems.

### v2

@@ Abstract
PAL Robotics, in Spain, builds a variety of Autonomous Mobile Robots (AMRs), which are deployed in diverse environments (e.g., warehouses, retail spaces, and offices), where they work alongside humans. Given that human behavior can be unpredictable and that AMRs may not have been trained to handle all possible unknown and uncertain behaviors, it is important to test AMRs under a wide range of human interactions to ensure their safe behavior. Moreover, testing in real environments with actual AMRs and humans is often costly, impractical, and potentially hazardous (e.g., it could result in human injury). To this end, we propose a Vision Language Model (VLM)-based testing approach (RVSG) for industrial AMRs developed together with PAL Robotics. Based on the functional and safety requirements, RVSG uses the VLM to generate diverse human behaviors that violate these requirements. We evaluated RVSG with several requirements and navigation routes in a simulator using the latest AMR from PAL Robotics. Our results show that, compared with the baseline, RVSG can effectively generate requirement-violating scenarios. Moreover, RVSG-generated scenarios increase variability in robot behavior, thereby helping reveal their uncertain behaviors.
@@ 2 Industrial Context
Within the RoboSAPIENS, self-adaptive behaviors are realized using the MAPLE-K loop [14], an extension of the well-established MAPE-K loop for self-adaptation in software systems. As presented in Fig. 1, in the MAPLE-K loop, an AMR continuously monitors (Monitor) its environment using dedicated sensors, analyzes the data (Analyze), and, upon detecting a situation that requires adaptation, plans (Plan), validates (Legitimate), and executes (Execute) the selected adaptation. Throughout this process, all information becomes part of the robot’s knowledge base (Knowledge), which can be accessed at any stage of the loop. The Legitimate phase is newly introduced in MAPLE-K.
@@ 3 Method
Fig. 2 presents the overview of RVSG, which aims to generate test scenarios that lead to the robot violating specific requirements. The workflow starts with RVSG Configuration (Section 3.1), where the F&S requirements, simulation world, and corresponding robot navigation routes are configured. Necessary test setups, such as test budget and stopping criteria, are also defined. After RVSG being configured, Environment Preprocessing (Section 3.2) begins, where Map Image Preprocessing is first performed to preprocess the simulation world map ( 1 in Fig. 2). Based on the preprocessed map, Environment Description Generation adopts a VLM to generate natural language descriptions of the environment by a three-turn prompting process at 2. After Environment Preprocessing being finished, the preprocessed maps, environment descriptions, together with the requirements and robot navigation routes, and other configurations are forwarded to the next phase for Test Scenario Generation.
Test Scenario Generation (Section 3.3) consists of three key components. The first is Human Configuration Generation by Multi-turn Conversation at 3, which generates the human configuration through multi-turn interaction with the VLM. Concretely, this component employs a Prompt Generator to construct the input prompts for the VLM. The Prompt Generator maintains a set of prompt templates, and can automatically select an appropriate template and instantiate a prompt based on the available information (e.g., preprocessed map images and requirements). Once the initial prompt is generated, the multi-turn conversation starts: each input–output pair forms a conversation turn and is returned to the Prompt Generator as the dialogue history for refining subsequent prompts. The process continues iteratively until all turns are completed, resulting in a final Human Configuration. Then the Human Configuration is simulated in Test Execution to test the robot navigation at 4. The feedback from test execution can be used by Prompt Generator to adjust the original prompts for improving effectiveness. Moreover, we design a Memory component to store all multi-turn conversations and test execution feedback collected throughout the experiments ( 5 in Fig. 2). The Prompt Generator leverages this accumulated experience to iteratively refine prompts, increasing diversity in scenario generation.
@@
Figure 2: Overview of RVSG during the iterative testing of autonomous mobile robots. A process decorated with Refer to caption involves a human operator; otherwise, it is fully automated.
3.1 RVSG Configuration
RVSG Configuration defines the global settings that can be specified in advance for each experiment, including the Requirement, Simulation World, Robot Navigation Route, and Test Setup. The Requirement (e.g., maintaining a safe distance between the robot and humans) is used to guide RVSG in generating test scenarios that violate it (e.g., collisions between the robot and humans). The Simulation World specifies the simulated environment for testing, supporting map acquisition ( 1 of Fig. 2), simulation execution ( 4), and feedback collection (details in Section 3.3.1). The Robot Navigation Route defines the AMR’s navigation path, including both a natural language description for VLM reasoning and parameter settings for simulation execution (e.g., a set of waypoints for robot navigation). The Test Setup contains various hyperparameter configurations related to the experiment, with further details provided in Section 4.2.
@@ 3.2.1 Map Image Preprocessing
3.2.2 Environment Description Generation
Leveraging the VLM’s strong visual reasoning, RVSG conducts multi-turn conversations to analyze step-by-step the labeled map and grid images, extracting layout and navigation information. As shown at 2 of Fig. 2, RVSG first provides the labeled map image and layout analysis instructions to the VLM to obtain a detailed analysis of environmental layout. Next, using the preserved dialogue history, RVSG sends the grid image and corresponding grid analysis instructions to the VLM, along with the map image that emphasizes the basic environmental information, and generates descriptions of the grid and navigation layouts. Finally, RVSG provides a summarization prompt to the VLM asking to summarize the key description for both map and grid images in JSON format. This output is then processed to extract and pair the corresponding descriptions for each image, storing them together as the current environment descriptions for subsequent testing and analysis. Given the potential for hallucinations or inaccuracies in the VLM’s output, the final environment descriptions are manually reviewed to ensure correctness.
@@ 3.3 Test Scenario Generation
3.3.1 Human Configuration Generation by Multi-turn Conversation
@@
For 3 in Fig. 2, RVSG utilizes Prompt Generator to automatically construct multi-turn conversation prompts based on predefined requirements and robot navigation routes, and preprocessed map images and environment descriptions. These prompts are then sequentially input into the VLM, enabling step-by-step reasoning through multi-turn conversations to produce intermediate outputs (e.g., scenario description) and final outputs (e.g., JSON format human configuration). Each conversation turn builds on previous context, ensuring the VLM continuously aligns with the test objective, understands the requirement, and generates the required output for each step. Moreover, by incorporating feedback from scenario simulations (at 4) and memory of previous scenarios (at 5), RVSG further optimizes configuration generation and enhances scenario diversity. Specifically, RVSG provides multiple prompt templates for five key reasoning steps: (1) Test Scenario Description Generation, (2) High-level Human Task Generation, (3) Low-level Human Waypoint Path Generation, (4) Human Configuration Validity Check and Revision, and (5) JSON Format Human Configuration Generation.
  • Test Scenario Description Generation analyzes the current environment and produces realistic textual descriptions of test scenarios involving human activities. These descriptions serve as the basis for subsequent human configuration.
  • High-level Human Task Generation allocates specific human tasks according to the generated scenario description, detailing concrete workflows and outputting the names of key labeled areas involved in these workflows.
  • Low-level Human Waypoint Path Generation builds on the high-level tasks, sequentially selecting valid waypoints from the preprocessed grid map based on the workflow and key areas to construct human waypoint paths. The high-level tasks and low-level waypoint paths together form the human configuration.
  • Human Configuration Validity Check and Revision checks the validity of the generated human configuration. Problematic configurations are automatically detected and revised by the VLM at this stage. For example, the module verifies whether the workflow aligns with the scenario description and whether the waypoint paths are feasible. If inconsistencies are found, the workflow is modified and the waypoint paths are updated accordingly.
  • JSON Format Human Configuration Generation provides the final human configuration in JSON format.
@@
Figure 3: Prompt templates, prompt field example, and scenario generation with feedback and memory in RVSG.
Fig. 3 (a) presents all seven prompt templates provided by RVSG. Each template focuses on one or two specific key reasoning step(s). Based on these templates, RVSG automatically fills in the provided information, such as requirements, into the corresponding fields to generate prompts for VLM reasoning and analysis. As shown in the Prompt Field Example in Fig. 3 (b), fields such as Environment Description, Robot Navigation Route, and Requirement are filled with content either preprocessed by the VLM or predefined in advance. Fields like Feedback Data and Historical Scenario Memory are dynamically added or updated during the RVSG execution, and are filled into the template in JSON format. Specifically, Feedback Data contains time-series information collected at each timestamp during simulation execution, including the positions of the robot and humans, as well as requirement violation metric values (e.g., the minimum distance between the robot and the humans). Historical Scenario Memory includes descriptions of previous scenarios, human configurations, and the corresponding violation metric values obtained in those scenarios. All templates and instantiated prompts are publicly available in our repository [45].
3.3.2 Test Execution
At 4 in Fig. 2, the Human Configuration is used to generate executable Human Agents in the simulator. To ensure realistic human behavior, RVSG employs HuNavSim [33], an agent controller based on behavior trees and the social force model, which closely simulates real-world human actions. The AMR under test is loaded with its designated navigation route and uses its navigation algorithm for environment perception and navigation decisions. Both human agents and the AMR perform their tasks in the simulator, enabling dynamic performance evaluation of the AMR along the navigation route, with simulation result data collected and computed throughout the process. The simulation results can be used to optimize the prompts at 3 for more effective scenario configuration generation. They are also stored in the Memory module (at 5) to enable continuous improvement of scenario generation. The simulation environment is consistent with that used in Environment Preprocessing, and real-time communication is maintained via Robot Operating System 2 (ROS2) [20].
3.3.3 Memory & Feedback
Memory module at 5 in Fig. 2 primarily stores all dialogue history from 3, along with the parsed scenario information and human configurations. It also retains the test execution feedback data from 4. This Memory module accumulates experience from scenario generation, enabling iterative optimization of the prompts in 3 to support the creation of more effective and diverse scenarios. By incorporating test execution feedback from the execution of each generated human configuration in the simulator, RVSG supports a feedback- and memory-based prompt optimization strategy for the improvement of test scenario generation, as illustrated in Fig. 3 (c).
@@ 3.3.3 Memory & Feedback
As shown in Fig. 3 (c), during the initial generation stage, RVSG selects a set of prompts, including the initial Test Scenario Description Generation prompt (Prompt1), High-level Human Task Generation prompt (Prompt3), Low-level Human Waypoint Path Generation prompt (Prompt4), Human Configuration Validity Check and Revision prompt (Prompt6), and JSON Format Human Configuration Generation prompt (Prompt7). These are combined with map images to form the initial multi-turn conversation prompts that guide the VLM in generating the initial human configuration, which is then executed in the simulator to obtain simulation results (e.g., the robot position and the requirement violation metric value at each timestamp). Subsequently, the feedback trigger determines whether to initiate feedback-based optimization. If optimization continues, RVSG combines the dialogue history with map images, Human Configuration Regeneration (Prompt5), Human Configuration Validity Check and Revision (Prompt6), and JSON Format Human Configuration Generation (Prompt7) to guide the VLM in refining the human configuration based on simulation feedback, without altering the initial scenario description. The optimized configuration is then executed in the simulator, after which new feedback is obtained. This feedback, including the requirement violation metric value, is used to evaluate the effectiveness of the optimized configuration and determine the degree of optimization. If further optimization is needed, the generated feedback can be used iteratively in subsequent optimization steps. If the feedback trigger indicates stopping feedback optimization (e.g., reaching a certain number of optimizations), RVSG then relies on the memory trigger to decide whether to use memory for generating more diverse scenarios. If memory-based optimization proceeds, RVSG replaces the initial Test Scenario Description Generation prompt (Prompt1) with the Diversity Test Scenario Description Generation prompt (Prompt2), keeping the other prompts unchanged to guide the VLM in creating the new test scenario. Moreover, in Prompt2, the Historical Scenario Memory field provides the descriptions of all previously generated test scenarios, along with their corresponding optimal configurations and requirement violation metric values, serving as references for generating more diverse scenarios. Otherwise, the process terminates, and all generated scenarios are output. All multi-turn conversation histories, simulation results, and corresponding scenario descriptions and human configurations are stored in the Memory module, supporting VLM-based scenario generation at different stages.
@@ 4.1 Subject System and Simulator
We test the latest logistics AMR robot, TIAGo OMNI Base from PAL Robotics [30]. It has four mecanum wheels for omnidirectional movement and two LiDAR scanners that provide a 360° unobstructed field of view, enabling it to detect and autonomously avoid obstacles to ensure safe navigation. To support simulation and testing, PAL Robotics offers an open-source TIAGO OMNI Base ROS 2 Simulation [29]. The simulator is built on Gazebo classic 11 [9] and provides a high-fidelity model of the TIAGo OMNI Base. The TIAGo OMNI Base adopts Nav2 [21] for both global navigation and local obstacle avoidance, while real-time communication is achieved through ROS 2 Humble [20]. Given the simulation platform and the ROS 2 version, we selected HuNavSim [33] as the default agent for human behavior control. It is capable of reacting to both the robot and humans. We used two human agents in our experiments. At the group level, HuNavSim uses the social force model [10] to simulate general crowd flow and avoidance behaviors in the presence of a robot, obstacles, or humans. At the individual level, it employs behavior trees to define and control more complex, personalized human actions. As a result, HuNavSim can provide realistic human behavior modeling, which benefits AMR navigation task testing and supports the development of navigation systems.